### v1

# Configurable photonic simulator for quantum field dynamics

Abstract
Quantum field simulators provide unique opportunities for investigating the dynamics of quantum fields through tabletop experiments. A primary drawback of standard encoding schemes is their rigidity: altering the theory, its coupling geometry, metric structure, or simulation time typically requires redesigning the experimental setup, which imposes strong constraints on the types of dynamics and theories that can be simulated. Here, we introduce the Optical Time Algorithm (OTA) as a unifying framework, enabling the efficient simulation of large classes of free quantum field dynamics using a single optical circuit design that separates the time from the Hamiltonian’s structure. By modifying the parameters of the optical elements, our method allows us to engineer timescales, coupling graphs, spacetime metrics, and boundary conditions, thereby facilitating the implementation of relativistic and non-relativistic, real- and complex-valued, short- and long-range quantum field theories on both flat and curved spacetimes. We exploit the OTA’s configurability to investigate the spreading of quantum correlations in space and time for theories with continuously varying coupling ranges. Relevant features predicted by quantum field theory can be observed on systems of $10-20$ modes, which paves the ground for experimental implementations.
I Introduction
40 years after Unruh’s proposal for mimicking Hawking radiation with sound waves [1], quantum field simulators have become a central tool for probing the non-equilibrium dynamics of quantum field theories [2, 3, 4, 5, 6, 7, 8]. By virtue of their high-level of control in the quantum regime, much attention has been drawn to ultracold atoms [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29], trapped-ions [30, 31, 32, 33, 34, 35, 36, 37, 38, 39] and Rydberg systems [40, 41, 42], all of which allow for probing few- and many-body physics. While each platform offers distinct opportunities, common approaches come with three main limitations: they operate far from the field theory limit, are often constrained by fixed coupling geometries, and allow for coherent dynamics only on short time scales. This poses significant challenges when attempting to simulate the dynamics of entire classes of proper field theories using a single setup.
With photonic devices on the rise [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55], simulating quantum fields using light is emerging as a compelling alternative [56, 57, 58, 59, 60, 61, 62]. In contrast to the majority of simulators leveraging spin degrees of freedom, photons with their bosonic nature are better suited for implementing scalar field theories – among which are theories paradigmatic in high-energy physics such as the relativistic scalar, long-range theories – being of particular relevance for quantum gravity [63] – and quantum fields in curved spacetime. This, together with the high configurability offered by integrated photonic circuits, raises the need for versatile yet efficient algorithms.
In this work, we propose the Optical Time Algorithm (OTA) as a unified framework for simulating relevant free scalar field theories on photonic platforms. In contrast to existing methods [64, 65, 66, 67, 68, 69] (see [70] for an overview), our decomposition into optical elements has a clear and physically intuitive structure: while a fixed array of beam splitters followed by a single layer of squeezers implements a largely arbitrary coupling geometry, the dynamics is entirely encoded in a single layer of phase shifters. This facilitates the simulation of the non-equilibrium dynamics of various quantum field theories without the need to ever alter the experimental design.
As an application of our method, we investigate information dynamics in terms of entanglement entropy and quantum mutual information [71, 72], also for non-local theories. We show that analytic predictions in the continuum for how quantum entanglement [73, 74, 75, 76] spreads in spacetime based on the quasi-particle picture [77, 78, 79, 80, 81, 82, 83] are in agreement with simulations with $10-20$ modes.
The remainder of this paper is structured as follows. We set the stage by recalling the necessary tools for describing continuous-variable quantum systems in Sec. II. We present our main result, the OTA, in Sec. III, and discuss its implications (we outsourced the proof to App. A). In Sec. IV, we showcase the OTA’s versatility by detailing how five rather different types of field theories can be simulated by minimally adjusting circuit parameters. As an application, we investigate the dynamics of quantum correlations after a quench in experimentally accessible regimes in Sec. V. We discuss our findings in Sec. VI.
Notation. We use natural units $\hbar=c=k_{\text{B}}=1$, denote quantum operators (classical variables) by bold (normal) letters $\bm{\rho}$ ($\gamma$), analogously for their traces $\textbf{Tr}\{\bm{\rho}\}$ ($\text{Tr}\{\gamma\}$) and work in $(1,d)$-dimensional spacetime with metric signature (+,,,)(+,-$\dots$,-)( + , - , … , - ).
II Preliminaries
We begin by mapping quantum fields to optical modes and introducing fundamental concepts in quantum optics [84, 85]. In particular, we introduce Gaussian operations and optical elements – phase shifters, beam splitters, and squeezers – serving as fundamental building blocks.
II.1 From quantum fields to discrete modes
We consider a real scalar quantum field $\phi(x)$ together with its conjugate momentum field $\pi(x)$ over a finite hypercube $x\in[0,L]^{d}$. We discretize spatial positions by introducing a hypercubic lattice with spacing $\epsilon>0$ and label discrete spatial modes with $j\in\mathcal{J}=\{1,\ldots,\tilde{N}\}^{d}$, such that $L=\tilde{N}\epsilon$, see Fig. 1(a). We choose both induced discretized fields to be dimensionless, i.e.,
$\phi(x)\rightarrow\phi_{j}=\phi(\epsilon j),\quad\pi(x)\rightarrow\pi_{j}=\epsilon^{d}\,\pi(\epsilon j),$ (1)
by means of a non-symmetric rescaling. This ensures that the corresponding quantum operators fulfill the standard canonical commutation relations (CCRs)
$[\bm{\phi}_{j},\bm{\pi}_{k}]=i\delta_{jk}\mathds{1},$ (2)
and agrees with quantum optics conventions. For simplicity, we relabel spatial indices to run from $1$ to $N$, where $N=\tilde{N}^{d}$ denotes the total number of modes.
We group [^1] Note that another prominent convention uses the grouping $\bm{r}=(\bm{\phi}_{1},\bm{\pi}_{1},\dots,\bm{\phi}_{N},\bm{\pi}_{N})^{\intercal}$ [84, 85]. the fields into a single vector $\bm{r}=(\bm{\phi}_{1},\dots,\bm{\phi}_{N},\bm{\pi}_{1},\dots,\bm{\pi}_{N})^{\intercal}$. Then, the CCRs (2) become
$[\bm{r},\bm{r}^{\intercal}]=i\Omega,$ (3)
where $\Omega$ is the so-called symplectic form,
$\Omega=\begin{pmatrix}\mathds{O}_{N}&\mathds{1}_{N}\\ -\mathds{1}_{N}&\mathds{O}_{N}\end{pmatrix}=(i\sigma_{y})\otimes\mathds{1}_{N},$ (4)
with $\sigma_{y}$ being the second Pauli matrix and $\mathds{O}_{N}$ ($\mathds{1}_{N}$) the $N\times N$ null (identity) matrix. The symplectic form is orthogonal and involutory up to a sign, $\Omega^{-1}=\Omega^{\intercal}=-\Omega$.
Figure 1: Pictorial representation of the map between scalar quantum fields (a) and the photonic simulation (b). By tuning the optical parameters, one can realize, for instance, arbitrary coupling geometries and timescales, allowing the probing of information dynamics in spacetime in general settings.
II.2 Gaussian unitaries and symplectic matrices
Free, i.e., non-interacting, field theories are characterized by second-order Hamiltonians, which result in Gaussian unitaries, as they map Gaussian states – states with Gaussian Wigner function – to Gaussian states. In this case, the most general Hamiltonian takes the form
$\bm{H}=\dfrac{1}{2}\bm{r}^{\intercal}H\bm{r}+\bm{r}^{\intercal}s,$ (5)
where $H$ denotes the symmetric and positive semi-definite Hamiltonian matrix and $s$ is a real $2N$-dimensional vector. Without loss of generality, we consider purely quadratic Hamiltonians and set $s\equiv 0$, since a linear term can be accounted for by a simple displacement.
The corresponding unitary time-evolution operator is
$\bm{U}(t)=e^{-it\bm{H}}=e^{-\frac{it}{2}\bm{r}^{\intercal}H\bm{r}}.$ (6)
The action of this unitary onto field operators can be equally described by symplectic matrices via $\bm{U}^{\dagger}\bm{r}\bm{U}=S\bm{r}$, where we introduced the symplectic transformation
$S(t)=e^{\Omega Ht}.$ (7)
In general, a matrix $S$ is called symplectic if it preserves the CCRs (3), i.e., $S\Omega S^{\intercal}=\Omega$, and the set of symplectic matrices forms the symplectic group $Sp(2N,\mathbb{R})$. Note also that they have unit determinant $\det S=1$.
A fundamental result for our analysis is the normal mode decomposition, also referred to as Williamson’s theorem [64]. Given a $2N\times 2N$ real and positive semi-definite matrix $H$ (e.g., the Hamiltonian matrix), there exists a symplectic diagonalizing (SD) transformation $S_{\text{SD}}\in Sp(2N,\mathbb{R})$ in the sense that [86, 87]
$S_{\text{SD}}^{\intercal}DS_{\text{SD}}=H\,\,\,\text{where}\,\,\,D=D_{N}\oplus D_{N},$ (8)
with the diagonal matrix
$D_{N}=\text{diag}(d_{1},\ldots,d_{N}),$ (9)
containing the so-called symplectic eigenvalues $d_{j}\in\mathbb{R}$. Importantly, the SD matrix is not unique [88, 89]: the product of two SD matrices satisfies $S_{\text{SD}}S_{\text{SD}}^{\prime-1}=G\in OrSp(2N)$, where
$\begin{split}OrSp(2N)&=\Big{\{}G=\begin{pmatrix}G_{1}&G_{2}\\ -G_{2}&G_{1}\end{pmatrix}:G_{1}^{\intercal}G_{2}=G_{2}^{\intercal}G_{1},\\ &\hskip 19.91684ptG_{1}^{\intercal}G_{1}+G_{2}^{\intercal}G_{2}=\mathds{1}_{N}\Big{\}}\end{split}$ (10)
defines the orthosymplectic group. It includes all transformations that are simultaneously orthogonal and symplectic, also known as passive transformations.
II.3 Optical elements
Finally, we introduce the fundamental optical elements that are relevant for a photonic simulation, see Fig. 1(b), which are conveniently grouped into passive (energy-conserving) and active (energy-changing) operations:
  1. (i)
    The (passive) single-mode phase shift (PS)
    $H_{\text{PS}}=\mathds{1}_{2},\quad S_{\text{PS}}(\varphi)=\begin{pmatrix}\cos{\varphi}&\sin{\varphi}\\ -\sin{\varphi}&\cos{\varphi}\end{pmatrix},$ (11)
    which implements a phase rotation $\varphi\in[0,2\pi).$
  2. (ii)
    The (passive) two-mode beam splitter (BS)
    $\begin{split}H_{\text{BS}}&=\sigma_{y}\otimes\sigma_{y},\\ S_{\text{\text{BS}}}(\theta)&=\begin{pmatrix}\cos{\theta}&\sin{\theta}&0&0\\ -\sin{\theta}&\cos{\theta}&0&0\\ 0&0&\cos{\theta}&\sin{\theta}\\ 0&0&-\sin{\theta}&\cos{\theta}\end{pmatrix},\end{split}$ (12)
    describing the interference of two modes with transmittivity $\cos^{2}(\theta)$ controlled by $\theta\in[-\pi,\pi]$ [^2] A physical beam splitter is described by $\theta\in[0,\pi/2]$. The remaining values can be realized with two additional $\pi$ phase shifts: one on the first mode before the BS, and one after the BS on mode two for $\theta\in(\pi/2,\pi]$, one before on mode two for $\theta\in[-\pi,-\pi/2)$, and one after on mode one when $\theta\in[-\pi/2,0)$..
  3. (iii)
    The (active) single-mode squeezer (Sq)
    $H_{\text{Sq}}=-\sigma_{x},\quad S_{\text{Sq}}(z)=\begin{pmatrix}e^{-z}&0\\ 0&e^{z}\end{pmatrix},$ (13)
    which decreases the uncertainty in one of the two fundamental fields at the cost of equally increasing the other, governed by $z\in\mathbb{R}$.
For multi-mode systems, we denote by $S_{\text{PS}}(\varphi)$ the $N$-mode phase shift with phase vector $\varphi=(\varphi_{1},\dots,\varphi_{N})^{\intercal}$ composed out of single-mode phase shifts $\varphi_{j}$; analogously for the $N$-mode squeezer. While the ordering of the individual single-mode components is irrelevant in these two cases, we fix the following triangular ordering for the $N$-mode BS (see Fig. 2)
$\begin{split}S_{\text{BS}}(\theta)&\equiv S_{\text{BS}}(\theta_{12})S_{\text{BS}}(\theta_{13})\dots S_{\text{BS}}(\theta_{1N})\\ &\hskip 39.26494pt\times S_{\text{BS}}(\theta_{23})\dots S_{\text{BS}}(\theta_{2N})\\ &\hskip 61.17325pt\times\dots\\ &\hskip 79.09862pt\times S_{\text{BS}}(\theta_{N-1N}),\end{split}$ (14)
with $S_{\text{BS}}(\theta_{jk})$ being the two-mode BS over modes $j,k$.
III Optical time algorithm
We are now ready to state and discuss our main result, the Optical Time Algorithm (OTA).
III.1 Main result
III.1.1 Isolating time dependencies
Building upon Williamson’s theorem, cf. Eq. (8), the symplectic time evolution matrix $S(t)$ corresponding to some Hamiltonian matrix $H$ can always be decomposed as (see App. A.1 for a proof)
$S(t)=S_{\text{SD}}^{-1}\,S_{\text{PS}}(\varphi(t))\,S_{\text{SD}}.$ (15)
Here, the phases $\varphi(t)=t\,(d_{1},\ldots,d_{N})^{\intercal}$ carry all time-dependencies in a linear fashion, and $d_{j}$ denote the symplectic eigenvalues of $H$. Thus, the dynamics is entirely encoded in a single layer of phase shifters.
III.1.2 OTA
For the vast majority of physically relevant scalar quantum field theories, the Hamiltonian matrix is of block-diagonal form $H=H^{\phi}\oplus H^{\pi}$ with commuting blocks $[H^{\phi},H^{\pi}]=\mathds{O}_{N}$. Under this assumption, we prove in App. A.2 that $S(t)$ admits the general decomposition
$S(t)=S_{\text{BS}}(\theta)\,S_{\text{Sq}}^{-1}(z)\,S_{\text{PS}}(\varphi(t))\,S_{\text{Sq}}(z)\,S_{\text{BS}}^{-1}(\theta).$ (16)
Remarkably, the design of the corresponding optical circuit, see Fig. 2, is independent of the theory under consideration. Instead, the structure of the Hamiltonian of interest is fully encoded in the parameters of the optical elements via
$\varphi(t)=t\,(d_{1},\ldots,d_{N})^{\intercal},$ (17a)
$z=-(\ln\gamma_{1},\ldots,\ln\gamma_{N})^{\intercal}/2,$ (17b)
$S_{\text{BS}}^{-1}(\theta)=P_{N}\oplus P_{N}.$ (17c)
Therein, $P_{N}=[p_{1},\ldots,p_{N}]^{\intercal}$ is composed of the real and orthonormal eigenvectors $p_{j}$ of $H^{\phi}$ (and $H^{\pi})$. The ratio of the corresponding eigenvalues $\lambda^{\phi}_{j}$ and $\lambda^{\pi}_{j}$ is related to the squeezing amplitudes
$\gamma_{j}^{2}=\lambda^{\phi}_{j}/\lambda^{\pi}_{j},$ (18)
while the phase shifts and symplectic eigenvalues $d_{j}$ of $H$ are given by
$d_{j}^{2}=\lambda^{\phi}_{j}\lambda^{\pi}_{j}.$ (19)
The triangular structure of the BS layers is dictated by Eq. (14), see the zoom-in in Fig. 2. It remains to find the BS parameters, for which we apply a standard iterative scheme, see App. A.2.3. Given a Hamiltonian matrix $H$, all circuit parameters can be calculated using [90].
Figure 2: Pictorial representation of the OTA for $N=5$.
The intuition behind the OTA is straightforward: the first beam-splitter layer diagonalizes the Hamiltonian matrix, thereby decoupling the modes. The following squeezer layer equalizes the prefactors of the fields $\phi$ and $\pi$, after which the (now independent) modes are evolved in time by phase shifters. Then, the second pair of a squeezer and beam-splitter layer transforms back to the original basis. As we will discuss in Sec. IV, the phase shift parameters $d_{j}$ often reduce to the dispersion relation $\omega_{j}$ of the field theory of interest. Further, we remark that the number of gates required by the OTA scales quadratically with the number of modes $N$, i.e.,
$n_{\text{G}}=n_{\text{Sq}}+n_{\text{PS}}+n_{\text{BS}}=N(N+2),$ (20)
consistent with other unitary decompositions in terms of single- and two-mode gates [91].
III.2 OTA with a single squeezer layer
III.2.1 General case
The obtained decomposition can be modified using known decompositions and adapted to specific experimental constraints. For instance, in order to reduce the number of squeezers, one can apply the Bloch-Messiah decomposition (see Ref. [70] for details) to the single-mode part of the OTA, leading to
$S(t)=S_{\text{BS}}(\theta)S_{\text{PS,1}}(t)S_{\text{Sq}}(t)S_{\text{PS,2}}(t)S_{\text{BS}}^{-1}(\theta).$ (21)
While the BS layers remain time-independent, both the squeezer and phase shifter arrays acquire time dependencies. This differs from the result one would obtain when directly applying the Bloch-Messiah decomposition, in which case, all gates are generally time-dependent.
What is more, the single-mode part can be summarized as so-called complex squeezers, defined by
$\begin{split}&H_{\text{CSq}}(\Phi)=-\sigma_{x}\cos\Phi-\sigma_{z}\sin\Phi,\\ &S_{\text{CSq}}(\zeta)=\cosh\xi\,\mathds{1}_{2}-\sinh\xi\begin{pmatrix}\cos\Phi&\sin\Phi\\ \sin\Phi&-\cos\Phi\end{pmatrix},\end{split}$ (22)
where $\zeta=\xi e^{i\Phi}$ with $\xi\in\mathbb{R}$ and $\Phi\in[0,\pi)$, which reduces to the standard squeezer, cf. Eq. (13), when $\Phi=0$, in which case $\zeta=\xi=z$. It is straightforward to check that
$S_{\text{Sq}}^{-1}(z)S_{\text{PS}}(\varphi)S_{\text{Sq}}(z)=S_{\text{CSq}}(\zeta),$ (23)
where the parameters on both sides are related via
$\begin{split}\cosh(2\xi)&=\cos^{2}(\varphi)+\sin^{2}(\varphi)\cosh(4z),\\ \tan\Phi&=-\dfrac{\cot\varphi}{\cosh(2z)}.\\ \end{split}$ (24)
Therefore, one can always replace the three inner single-mode layers of the OTA (16) (see also Fig. 2) with a single layer of complex squeezers. As was the case above, the price to pay is that the complex squeezer parameters $\xi$ and $\Phi$ are both time-dependent, since $\varphi_{j}(t)=d_{j}t$.
III.2.2 Coherent inputs
So far, we have worked at the level of time-evolution operators without assuming a specific input state. Interestingly, the OTA requires fewer gate layers when considering the standard resource in quantum optics experiments, that is, coherent input states. We switch to mode operators $\bm{a}_{j}=(\bm{\phi}_{j}+i\bm{\pi}_{j})/\sqrt{2}$ to define coherent states as displaced vacuum states |$\ket{\alpha}=\bm{D}(\alpha)\ket{0}$, where $\bm{D}(\alpha)=e^{\alpha\bm{a}^{\dagger}-\alpha^{*}\bm{a}}$ with $\bm{a}=(\bm{a}_{1},\dots,\bm{a}_{N})^{\intercal}$ (analogously for the vector of complex phases $\alpha\in\mathbb{C}^{N}$). Here, the vacuum state |$\ket{0}$ is the product state annihilated by all $\bm{a}_{j}$.
Noting that passive transformations map coherent states onto coherent states [84], the first BS layer merely modifies the coherent states’ amplitudes $\alpha\to\alpha^{\prime}$, i.e., the first moment of the state. Remarkably, thanks to the simplifications involving the complex squeezers in Eq. (23), the OTA can be rearranged to a layer of time-dependent squeezers and a single time-independent BS array, to wit
$S(t)\bra{\alpha}\bm{r}\ket{\alpha}=S_{\text{BS}}(\theta)S_{\text{CSq}}(\zeta(t))\bra{\alpha^{\prime}}\bm{r}\ket{\alpha^{\prime}}.$ (25)
We sketch the corresponding circuit in Fig. 3.
For the vacuum input, both coherent amplitudes vanish, i.e., |$\ket{\alpha^{\prime}}=\ket{\alpha}=\ket{0}$. In this case, the OTA reduces to processing a single-mode squeezed vacuum state input with passive linear optics, as realized in Gaussian boson samplers [92, 93, 51, 54]. Note that in the standard formulation, the Gaussian boson sampler includes particle-number-resolving detectors at the output [94, 95, 96], while no particular readout is preferred by the OTA. Instead, for quantum field simulation, the detection method can be chosen to suit the application of interest.
Figure 3: For coherent inputs, the OTA reduces to a complex squeezer layer followed by passive linear gates. The adjusted coherent inputs $\alpha^{\prime}$ and the complex-valued squeezing parameters $\zeta=\xi e^{i\Phi}$ are time-dependent, while the beam splitters remain time-independent.
IV Simulating quantum fields
Next, we discuss some representative classes of quantum field theories that the OTA can efficiently simulate. For simplicity, we consider one spatial dimension $d=1$ and periodic boundary conditions. Generalizations to higher dimensions are straightforward: the total mode number scales as $\tilde{N}^{d}$, affecting only the structure of the BS layers. Similarly, boundary conditions are arbitrary.
IV.1 General Hamiltonian structure
It is a central observation that many scalar field Hamiltonians become diagonal in momentum space. In the discretized picture, this is intimately related to the $\phi$-block of the Hamiltonian matrix $H$ corresponding to a circulant (and by definition, real and symmetric) matrix
$\begin{split}H^{\phi}&=\text{circ}(h^{\phi}_{1},h^{\phi}_{2},\dots,h^{\phi}_{N})\\ &=\begin{pmatrix}h^{\phi}_{1}&h^{\phi}_{N}&\dots&h^{\phi}_{3}&h^{\phi}_{2}\\ h^{\phi}_{2}&h^{\phi}_{1}&h^{\phi}_{N}&\dots&h^{\phi}_{3}\\ \vdots&h^{\phi}_{2}&h^{\phi}_{1}&\ddots&\vdots\\ h^{\phi}_{N-1}&\dots&\ddots&\ddots&h^{\phi}_{N}\\ h^{\phi}_{N}&h^{\phi}_{N-1}&\dots&h^{\phi}_{2}&h^{\phi}_{1}\end{pmatrix},\end{split}$ (26)
with the symmetric constraint $h_{j}=h_{N+2-j}$ understood (such that it has only $\lfloor N/2\rfloor+1$ independent elements).
Indeed, circulant matrices are diagonalized by a discrete Fourier transform (DFT). Hence, the required parameters for the phase shifters and squeezers, cf. Eqs. (17a) and (17b), respectively, can be computed analytically in full generality. The real eigenvalues $\lambda_{j}^{\phi}$ of $H^{\phi}$ read
$\lambda_{j}^{\phi}=\sum_{k=1}^{N}h^{\phi}_{k}\cos\dfrac{2\pi(j-1)(k-1)}{N},$ (27)
and typically encode the dispersion of the theory $\lambda_{j}^{\phi}\propto\omega_{j}$.
To construct the corresponding set of real eigenvectors $P_{N}$, we consider trigonometric functions instead of the usual complex-valued exponentials. The first eigenvalue $\lambda_{1}^{\phi}=\sum_{k=1}^{N}h_{k}$ has the constant eigenvector $p_{1}=\frac{1}{\sqrt{N}}\left[1,\dots,1\right]^{\intercal}$. The eigenvalues in $2\leq j\leq\lfloor N/2\rfloor$ are two-fold degenerate with respect to reflection at $N/2+1$, i.e., $\lambda_{j}^{\phi}=\lambda_{N+2-j}^{\phi}$. Each pair of such eigenvalues comes with two real eigenvectors
$\begin{split}(p_{j}^{\text{cos}})_{k}&=\tfrac{2}{\sqrt{N}}\cos\tfrac{2\pi(j-1)(k-1)}{N},\\ (p_{j}^{\text{sin}})_{k}&=\tfrac{2}{\sqrt{N}}\sin\tfrac{2\pi(j-1)(k-1)}{N}.\end{split}$ (28)
For even $N$, there is an additional eigenvalue $\lambda_{N/2+1}=\sum_{k=1}^{N}h^{\phi}_{k}(-1)^{k-1}$ corresponding to the alternating eigenvector $p_{N/2+1}=\frac{1}{\sqrt{N}}\left[1,-1,\dots\right]^{\intercal}$. We emphasize that the eigenvectors are independent of the matrix elements $h_{j}^{\phi}$; hence, the BS layers are identical for all Hamiltonians diagonalized by a DFT.
IV.2 Relativistic theory
The archetypal scalar field theory, that is, the relativistic, real, and massive scalar field, is characterized by the Lagrangian density
$\mathcal{L}_{\text{R}}=\dfrac{1}{2}\left[(\partial_{t}\phi)^{2}-(\partial_{x}\phi)^{2}-m^{2}\phi^{2}\right],$ (29)
where $m\geq 0$ denotes the mass. A special case is the massless limit $m\to 0$, which renders the theory conformal, and serves as an effective theory for various phenomena in the low-energy regime: phononic excitations of a Bose-Einstein condensate (BEC) [97]; the Tomonaga-Luttinger-Liquid after bosonization [98] describing linear excitations of interacting fermions in, e.g., quantum wires [99]; excitations of the XXZ Heisenberg spin chain at criticality [100]. The massive variant appears, e.g., for the linear fluctuations between two strongly tunnel-coupled BECs [28, 29].
After identifying the conjugate momentum field $\pi=\partial_{t}\phi$ and discretizing the theory as discussed in Sec. II, the Hamiltonian to be simulated reads
$\bm{H}_{\text{R}}=\dfrac{1}{2\epsilon}\sum_{j\in\mathcal{J}}\left[\bm{\pi}^{2}_{j}+(\bm{\phi}_{j+1}-\bm{\phi}_{j})^{2}+\epsilon^{2}m^{2}\bm{\phi}^{2}_{j}\right],$ (30)
with $\epsilon>0$ denoting the lattice spacing and $j\in\mathcal{J}=\{1,\ldots,N\}$ being a discrete label, such that $L=N\epsilon$.
It is evident that this Hamiltonian is of the form required for applying the OTA since the corresponding Hamiltonian matrix decomposes as $H_{\text{R}}=H_{\text{R}}^{\phi}\oplus H_{\text{R}}^{\pi}$ and the diagonal matrix $H_{\text{R}}^{\pi}=\mathds{1}_{N}/\epsilon$ commutes with any other matrix. Further, the $\phi$-block is indeed circulant as
$H_{\text{R}}^{\phi}=\begin{pmatrix}\mu&\nu&\ldots&0&\nu\\ \nu&\mu&\ldots&0&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ 0&0&\ldots&\mu&\nu\\ \nu&0&\ldots&\nu&\mu\end{pmatrix},$ (31)
where we introduced $\mu=\epsilon m^{2}-2\nu$ and $\nu=-1/\epsilon$.
Following Sec. IV.1, $H_{\text{R}}^{\phi}$ is diagonalized by a DFT, and hence the phase shifter and squeezer parameters follow directly from (27) and are given by $d_{j}=\omega_{j},\gamma_{j}=\epsilon\,\omega_{j}$ with the discretized relativistic dispersion
$\omega^{2}_{j}=m^{2}+\dfrac{4}{\epsilon^{2}}\sin^{2}\left[\tfrac{\pi(j-1)}{N}\right].$ (32)
In App. B, we exemplarily list all parameters for $N=5$. We remark that the massless limit $m\to 0$ causes the first pair of parameters to vanish $d_{1}=\gamma_{1}=0$, independent of $\epsilon$ and $N$. At first glance, this would require infinite squeezing $z_{1}\to\infty$ in the first mode. However, since the corresponding phase shift disappears as well, $\varphi_{1}(t)=0$, the two squeezer transformations cancel each other. Hence, any non-zero value can be chosen for $z_{1}$ without affecting the physics.
IV.3 Non-local theories
With the relativistic scalar serving as our baseline, we investigate how far the OTA can stretch by altering relevant physical properties. A key aspect of a field theory is its coupling range. Many physical processes are inherently local in the sense that changing a physical quantity at a specific spatial position only affects its values in the vicinity of this point. In the Lagrangian, locality manifests as two spatial derivatives acting on the field in the continuum $\sim(\partial_{x}\phi)^{2}$ and nearest-neighbor coupling terms in the discrete setting $\sim(\phi_{j+1}-\phi_{j})^{2}$, as is the case for the relativistic theory, cf. Eqs. (29) and (30), respectively.
In the following, we consider non-local theories, in particular, the class of long-range theories (see [101] for a review), characterized by a power-law decay of couplings at large distances. The long-range dynamics of both non-interacting and interacting systems have been extensively studied in trapped-ion [40, 31, 32, 34, 35, 36, 38, 39] and Rydberg [41] setups as well as optical cavities [24], with varying control over the coupling range: Dipolar interactions are characteristic for Rydberg atoms and magnetic atoms or molecules, while trapped-ion systems can also realize all-to-all couplings.
To study long-range effects beyond the many-body regime, i.e., in a proper field-theoretic setting using the OTA, we consider the family of theories where the Laplace operator is raised to some positive power $\alpha>0$, thereby controlling the range of the couplings. The Lagrangian reads
$\mathcal{L}_{\text{FL}}=\dfrac{1}{2}\left[(\partial_{t}\phi)^{2}-\phi\,(-\partial^{2}_{x})^{\alpha/2}\phi-m^{\alpha}\phi^{2}\right],$ (33)
and is referred to as fractional Laplacian theory [102, 103, 104, 105, 106, 107, 108]. The discrete Hamiltonian evaluates to
$\bm{H}_{\text{FL}}=\dfrac{1}{2\epsilon}\sum_{j\in\mathcal{J}}\Big{[}\bm{\pi}^{2}_{j}+\sum_{j^{\prime}\in\mathcal{J}}\bm{\phi}_{j}f_{jj^{\prime}}(\alpha)\bm{\phi}_{j^{\prime}}+\epsilon^{2}m^{\alpha}\bm{\phi}^{2}_{j}\Big{]},$ (34)
with $f_{jj^{\prime}}$ being a circulant and symmetric matrix, which, in the infinite volume limit, is given by
$f_{jj^{\prime}}(\alpha)=-\dfrac{\Gamma(-\tfrac{\alpha}{2}+j-j^{\prime})\Gamma(\alpha+1)}{\pi\Gamma(1+\tfrac{\alpha}{2}+j-j^{\prime})}\sin\left(\dfrac{\alpha\pi}{2}\right).$ (35)
We refer to App. C for the corresponding expression on a finite-size lattice.
This model captures all essential features of non-local theories: The local couplings associated with the relativistic theory (29) are obtained for $\alpha=2$, in which case $f_{jj^{\prime}}(2)=2\delta_{jj^{\prime}}-\delta_{jj^{\prime}+1}-\delta_{jj^{\prime}+N}$. When $\alpha<2$, the theory exhibits long-range couplings up to all-to-all couplings in the limit $\alpha\to 0$. Additionally, $\alpha>2$ describes ultra-short-range couplings constituting higher-derivative theories that are often considered in the context of quantum gravity [109, 63].
The Hamiltonian matrix decomposes $H_{\text{FL}}=H_{\text{FL}}^{\phi}\oplus H_{\text{FL}}^{\pi}$ with circulant and symmetric $(H_{\text{FL}}^{\phi})_{jj^{\prime}}=\epsilon m^{\alpha}\delta_{jj^{\prime}}+f_{jj^{\prime}}(\alpha)$ and $H_{\text{FL}}^{\pi}=\mathds{1}_{N}/\epsilon$. Hence, the OTA’s parameters are again given by $d_{j}=\omega_{j}(\alpha),\gamma_{j}=\epsilon\,\omega_{j}(\alpha)$, with the fractional dispersion
$\omega^{2}_{j}(\alpha)=m^{\alpha}+\dfrac{2^{\alpha}}{\epsilon^{2}}\left\lvert\sin^{\alpha}\left[\tfrac{\pi(j-1)}{N}\right]\right\rvert.$ (36)
We note that other types of non-local theories can also be simulated by the OTA, such as the exponentiated fractional Laplacian considered in Refs. [110, 105], for which the $\pi$-component of the Hamiltonian matrix is also proportional to the identity.
IV.4 Complex fields and the free Bose gas
When generalizing to complex-valued fields, the relativistic Lagrangian density attains the form
$\mathcal{L}_{\text{C}}=(\partial_{t}\psi^{*})(\partial_{t}\psi)-(\partial_{x}\psi^{*})(\partial_{x}\psi)-m^{2}\psi^{*}\psi.$ (37)
This model describes particles and anti-particles with their particle number difference being a conserved charge corresponding to a $U(1)$ symmetry. After decomposing the complex field into two real-valued fields $\phi_{1}$ and $\phi_{2}$ via $\psi=(\phi_{1}+i\phi_{2})/\sqrt{2}$, the Lagrangian decouples $\mathcal{L}_{\text{C}}=\mathcal{L}_{\text{R},1}+\mathcal{L}_{\text{R},2}$. Hence, the complex relativistic theory can be quantum-simulated by doubling the number of modes.
Of particular interest is the non-relativistic limit in Eq. (37). To this end, we factor out fast-oscillating contributions by rescaling the field $\psi\to e^{-i(m-V)t}\psi/\sqrt{2m}$, where $V>0$ represents an external field. Then, neglecting second-order time derivatives, which amounts to the limit $c\to\infty$ after reinstating units, gives
$\mathcal{L}_{\text{NRC}}=i\psi^{*}\partial_{t}\psi-\dfrac{1}{2m}(\partial_{x}\psi^{*})(\partial_{x}\psi)-V\psi^{*}\psi,$ (38)
which describes, for instance, a free Bose gas in an external trapping potential $V=V(x)$ [97].
For discretizing the non-relativistic theory, we parametrize the complex field in terms of the two canonical fields as $\bm{\psi}=(\bm{\phi}+i\bm{\pi})/\sqrt{2\epsilon}$, with the additional factor $1/\sqrt{\epsilon}$ guaranteeing that (2) is fulfilled. This results in the decoupled Hamiltonian
$\begin{split}\bm{H}_{\text{NR}}=\frac{1}{2}\sum_{j\in\mathcal{J}}\Big{\{}&\frac{1}{2m\epsilon^{2}}\left[(\bm{\phi}_{j+1}-\bm{\phi}_{j})^{2}+(\bm{\pi}_{j+1}-\bm{\pi}_{j})^{2}\right]\\ &+V_{j}\left(\bm{\phi}_{j}^{2}+\bm{\pi}_{j}^{2}\right)\Big{\}}.\end{split}$ (39)
After inverting $\bm{a}_{j}=(\bm{\phi}_{j}+i\bm{\pi}_{j})/\sqrt{2}$, the latter attains the form of a non-interacting Bose-Hubbard model [111, 112]
$\begin{split}\bm{H}_{\text{BH}}=\frac{1}{2}\sum_{j\in\mathcal{J}}\Big{[}&-\dfrac{1}{m\epsilon^{2}}\left(\bm{a}^{\dagger}_{j+1}\bm{a}_{j}+\bm{a}^{\dagger}_{j}\bm{a}_{j+1}\right)\\ &+2\Big{(}V_{j}+\dfrac{1}{m\epsilon^{2}}\Big{)}\bm{a}^{\dagger}_{j}\bm{a}_{j}\Big{]}.\end{split}$ (40)
describing ultracold atoms hopping between neighboring sites in an optical lattice, with hopping amplitude $-1/(m\epsilon^{2})$ and a site-dependent potential $2[V_{j}+1/(m\epsilon^{2})]$. The corresponding Hamiltonian matrix decomposes as $H_{\text{NR}}=H^{\phi}_{\text{NR}}\oplus H^{\pi}_{\text{NR}}$, with $H^{\phi}_{\text{NR}}=H^{\pi}_{\text{NR}}$. Since the two blocks agree, we find $\gamma_{j}=1$, such that no squeezing is required. Thus, the non-relativistic theory, or equivalently, the Bose-Hubbard model, can be simulated with passive linear optics only.
A general trapping profile $V_{j}$ breaks the circulanticity of both blocks of $H$, which read
$H_{\text{NR}}^{\phi}=H_{\text{NR}}^{\pi}=\begin{pmatrix}\mu_{1}&\nu&\dots&0&\nu\\ \nu&\mu_{2}&\nu&\dots&0\\ \vdots&\nu&\mu_{3}&\ddots&\vdots\\ 0&\dots&\ddots&\ddots&\nu\\ \nu&0&\dots&\nu&\mu_{N}\end{pmatrix},$ (41)
with $\mu_{j}=V_{j}-2\nu$ and $\nu=-1/(2m\epsilon^{2})$. There exists no closed-form analytic solution for the eigenvalues and eigenvectors of the latter matrix in general; hence, numerical methods must be used to determine the circuit parameters. When the potential is constant $V\equiv V_{j}$, the parameter $\mu\equiv\mu_{j}=V-2\nu$ amounts to the chemical potential, and the blocks of the Hamiltonian matrix reduce to the relativistic form (31). In this case, we obtain $d_{j}=\omega_{j}$ for the phase shifts, with the non-relativistic dispersion
$\omega_{j}=V+\frac{2}{m\epsilon^{2}}\sin^{2}\left[\tfrac{\pi(j-1)}{N}\right].$ (42)
IV.5 Curved spacetimes
Since more than two decades, simulating quantum fields in curved spacetime [113, 114] is of substantial interest to experimentally probe phenomena that remain inaccessible in the nightsky, including cosmological particle production [10, 19, 61, 25, 26, 27], the Hawking [9, 56, 12, 21, 115] and Unruh [22] effects, and rotating analog black holes [116, 117]. The standard approach in the field of analog gravity is to consider low-energy perturbations in a fluid, whose effective equation of motion is that of a massless Klein-Gordon field in a curved spacetime dictated by the background fluid flow in the form of the so-called acoustic metric [118]. It is well-known, however, that this one-to-one correspondence breaks down for $1+1$ spacetime dimensions as conformal symmetry prevents the wave equation of the perturbations from being rewritten in terms of a proper metric – thereby restricting many experiments to qualitative analogies. Furthermore, simulating a massive field typically requires two-component fluids [119]. In the following, we demonstrate that a photonic simulation with the OTA overcomes both of these problems.
We start again from the relativistic theory (29) and generalize to arbitrary spacetime geometries by introducing a metric $g_{\mu\nu}$ in the Lagrangian density,
$\mathcal{L}_{\text{CS}}=\dfrac{1}{2}\sqrt{-g}\left[g^{\mu\nu}(\partial_{\mu}\phi)(\partial_{\nu}\phi)-m^{2}\phi^{2}\right],$ (43)
where $\mu,\nu=\{t,x\}$ label spacetime coordinates (with the Einstein convention understood) and $\sqrt{-g}=\sqrt{-\text{det}(g_{\mu\nu})}$ being the invariant volume element. In two dimensions, every metric is conformally flat and thus proportional to the standard Minkoskwi metric up to a conformal factor, i.e., $g_{\mu\nu}=f^{2}(t,x)\eta_{\mu\nu}=f^{2}(t,x)\text{diag}(1,-1)$. Prominent choices for $f(t,x)$ include static situations, e.g., two-dimensional Anti-de-Sitter space AdS2, for which $f(x)=1/x^{2}$, or, scenarios where a causal horizon arises for the quasi-particle excitations, such as, for instance, $f(x)\propto 1+\tanh(x)$ [21, 115]. On the other hand, there are purely time-dependent settings: choosing an FLRW-type metric $g_{\mu\nu}=\text{diag}[1,-a^{2}(t)]$ encodes cosmological expansion of a homogeneous and isotropic universe [10, 25, 26, 27].
By exploiting conformal flatness, the corresponding discretized Hamiltonian attains the simple form
$\bm{H}_{\text{CS}}=\dfrac{1}{2\epsilon}\sum_{j\in\mathcal{J}}\bigg{[}\bm{\pi}_{j}^{2}+(\bm{\phi}_{j+1}-\bm{\phi}_{j})^{2}+\epsilon^{2}m_{\text{eff}}^{2}(t,\epsilon j)\bm{\phi}_{j}^{2}\bigg{]},$ (44)
with a spacetime-dependent effective mass $m_{\text{eff}}(t,x)=mf(t,x)$ (for the FLRW metric, we find $f(t)=a(t)$ and the Hamiltonian must be divided by $a(t)$). As was the case for Minkowski spacetime, the Hamiltonian (44) is block diagonal $H_{\text{CS}}=H_{\text{CS}}^{\phi}\oplus H_{\text{CS}}^{\pi}$ with $H_{\text{CS}}^{\pi}=\mathds{1}_{N}/\epsilon$. However, analogous to the spatially varying trapping profile in the non-interacting Bose-Hubbard model [see Eq. (40)], spatial curvature breaks the circulanticity of the $\phi$-block, which is of the form (41) with $\mu_{j}(t)=\epsilon m^{2}_{\text{eff}}(t,\epsilon j)-2\nu$ and $\nu=-1/\epsilon$.
In the absence of spatial curvature, the latter reduces to the relativistic expression (31) with a time-dependent effective mass, which can still be diagonalized by means of the same DFT. More precisely, the fixed form of the DFT eigenvectors (28) guarantees the BS layer to remain unchanged, while the eigenvalues, i.e., the dispersion, changes over time $\omega_{j}=\omega_{j}(t)$, which implies non-linear time-dependencies in the phases $\varphi_{j}(t)$ and the squeezing parameters $z_{j}(t)$. For the FLRW universe, the overall factor $1/a(t)$ can be reabsorbed by mapping the physical time $t$ to $\int_{0}^{t}\mathrm{d}t^{\prime}a(t^{\prime})$, which is reminiscent of the conformal time in cosmology [120]. In static situations, the OTA maintains its form, with all time dependencies isolated in a single phase shift layer, while the circuit parameters must be computed numerically in general.
V Information dynamics
As an application, we study the spreading of entanglement in spacetime after a global quantum quench for varying coupling range. We focus on identifying a regime where continuum features emerge while maintaining the number of modes at an experimentally feasible level.
V.1 Quench dynamics
We consider the experimentally friendly vacuum state |$\ket{0}$ as the input state, in which case the OTA reduces to processing squeezed vacuum inputs with linear optics, as shown in Fig. 3. The input state describes zero photons, which is the ground state of the pre-quench Hamiltonian
$\bm{H}_{0}=\dfrac{1}{2\epsilon}\sum_{j\in\mathcal{J}}(\bm{\pi}^{2}_{j}+\bm{\phi}^{2}_{j}).$ (45)
Importantly, |$\ket{0}$ belongs to the class of Gaussian states, which are characterized by their mean and covariance
$r=\textbf{Tr}\{\bm{\rho}\bm{r}\},\quad\sigma_{jj^{\prime}}=\dfrac{1}{2}\textbf{Tr}\left\{\bm{\rho}\{\bm{r}_{j}-r_{j},\bm{r}_{j^{\prime}}-r_{j^{\prime}}\}\right\},$ (46)
respectively, with the latter encoding all two-point correlation functions. The ground state of (45) is Gaussian with $r_{0}=0$ and $\sigma_{0}=(1/2)\mathds{1}_{2N}$.
We time-evolve the state under the relativistic theory (30), which constitutes a quantum quench of both the mass and the couplings at $t=0$. We remark that the often-considered mass quench would instead require the input state to be the ground state of the Hamiltonian (30) with a different mass – thus describing no quasi-particles (with respect to the initial theory) instead. In the relativistic case, this would necessitate additional squeezers during state preparation, see e.g., Ref. [59].
During time evolution, the state remains Gaussian, with $r(t)=S(t)r_{0}=0$ and $\sigma(t)=S(t)\sigma S(t)^{\intercal}$ [84]. This greatly simplifies the study of quantum entanglement and correlations, as standard quantum information measures can be expressed in terms of the covariance matrix only. From an experimental point of view, the covariance matrix can be reconstructed through homodyne measurements [121], with the number of measurements being proportional to (sub-)system size. Multimode homodyne measurements, also relevant in the context of continuous-variable quantum computing and cluster state generation [122], have been demonstrated on various experimental platforms [123, 124, 125, 55, 54].
V.2 Benchmarks: relativistic fields
V.2.1 Entanglement generation
We consider a spatial subregion $[0,\ell]$ with $\ell\leq L$, see inset of Fig. 4(a). As the modes are coupled, the degrees of freedom in $[0,\ell]$ become entangled with the ones in the complementary region $(\ell,L]$. Thus, while the global state $\bm{\rho}(t)$ remains pure during the unitary evolution, the local state $\bm{\rho}(t,l)=\textbf{Tr}_{L-\ell}\{\bm{\rho}(t)\}$ follows the evolution of an open system described by a completely positive trace-preserving map [126]. We assess entanglement generation via the Rényi-2 entanglement entropy, which, for Gaussian states, attains the simple form
$S_{2}(t,\ell)=-\ln\textbf{Tr}\{\bm{\rho}^{2}(t,\ell)\}=\dfrac{1}{2}\ln\det[2\sigma(t,\ell)],$ (47)
where $\sigma(t,\ell)$ denotes the covariance associated with the local state $\bm{\rho}(t,\ell)$. Further, we remark that, for Gaussian states, the latter is equivalent to the recently introduced subtracted Wigner entropy [127, 128, 129].
Figure 4: (a) Time evolution of the Rényi-2 entanglement entropy (47) (scaled by $10^{2}$) of the subregion $[0,\ell]$ following a quench from the optical vacuum (45) to the relativistic theory (30) with $\ell=L/5,m=1,\epsilon=2$ for varying $L=N\epsilon$ against the field-theoretic prediction (48) (black dashed curve). The light-cone-like motion of quasi-particles generates entanglement (see inset), leading to a linear increase in entanglement entropy until the time $\tau/N=\ell/(2Nv_{\text{max}})\approx 0.48$ (gray box) that the fastest quasi-particle pairs require to end up in distinct regions. (b) Same analysis for the Rényi-2 mutual information (49) between two spatial regions separated by a distance $d=L/5$. Correlations can only emerge after $t\geq\tau$, i.e., when the fastest quasi-particles have bridged the distance $d$.
Our main interest is to benchmark our quantum field simulator against QFT predictions. In the field theory context, the non-equilibrium dynamics of integrable models is understood in terms of a quasi-particle description [77, 78, 79, 80, 81, 82, 83] [see inset in Fig. 4a)]. Within this framework, the input state at $t=0$ acts as a source of quasi-particle pairs – momentum modes of opposite momenta $p$ and $-p$ – which ballistically spread correlations throughout the system after the quench. Then, at any given time $t>0$, a pair contributes to the entanglement entropy if one particle ends up in the subregion of interest, while the other resides in the complementary region. Combining this with the fact that the non-equilibrium steady-state for $t\to\infty$ attains the form of a generalized Gibbs ensemble (GGE) locally, i.e., $\bm{\rho}(\infty,l)\propto\text{exp}(-\sum_{p}y_{p}\bm{n}_{p})$, where $\bm{n}_{p}=\tilde{\bm{a}}^{\dagger}_{p}\tilde{\bm{a}}_{p}$ counts (conserved) quasi-particles with momenta $p$ and $y_{p}$ are Lagrange multipliers [130, 131], leads to the QFT prediction in the infinite volume limit
$S_{2}(t,\ell)=\int\dfrac{\mathrm{d}p}{2\pi}s_{2}(p)\min(2|v(p)|t,\ell).$ (48)
Therein, $v(p)=\mathrm{d}\omega(p)/\mathrm{d}p=\sin(\epsilon p)/(\epsilon\omega(p))$ denotes the group velocity of the bosonic excitations and $s_{2}(p)$ is the thermodynamic Rényi-2 entropy density associated with the GGE. We refer to App. D.1 for finite-size corrections, to App. D.2 for an analytic calculation of $s_{2}(p)$, and to App. D.3 for a comment on divergences appearing in the continuum limit $\epsilon\to 0$.
We compare the photonic simulation of the quench dynamics with the QFT prediction (48) (black dashed curve) in Fig. 4(a) for $\ell=L/5,m=1,\epsilon=2$ and varying system sizes $N=5-25$ (see App. B for circuit parameters when $N=5$). Shortly after the quench, we observe a linear growth of the Rényi-$2$ entanglement entropy $S_{2}(t,l)\propto t$ up to $t\leq\tau=\ell/(2v_{\text{max}})$, see gray region. The time $\tau$ marks the moment where two quasi-particles moving towards each other with the maximum velocity $v_{\text{max}}$ allowed by the relativistic dispersion have traveled a distance $\ell$ (see overlapping light cones in the inset).
After $t\geq\tau$, also the slower quasi-particles start to contribute to the entanglement between the subregion $[0,\ell]$ and its complement, but the growth of the entanglement entropy slows down due to saturation effects. For an infinitely extended system, the entropy would settle to its stationary value dictated by the GGE. For a finite system with periodic boundaries, however, the quasi-particles return to their initial positions every time they have traversed the system. This leads to so-called revivals in the entanglement entropy, marked by a temporary dip around the time $L/(2v_{\text{max}})=5\tau$ the fastest quasi-particles took to cross the system once. As the quasi-particles’ momenta vary, the entanglement is never entirely destroyed, and the dip attains a finite width.
Over the whole time range considered, the photonic simulation is in remarkable agreement with the QFT prediction – even for small system sizes $N\sim 10$ [51, 54]. Both the initial linear scaling and the revival dip(s) are particularly robust features that are clearly visible for even smaller mode numbers. Importantly, the circuit parameters fall into experimentally accessible regimes, see App. B. In general, implementing strong squeezing is most challenging. Here, the maximum squeezing required $z_{\text{max}}$, which corresponds to the zero-momentum mode, is independent of system size and evaluates to $\lvert z_{\text{max}}\rvert\approx 0.5\approx 4.3$dB – far below current records around $15$dB [132]. This demonstrates that the photonic simulation of entanglement dynamics in quantum fields is well within the reach of current experimental technologies.
Figure 5: Spatiotemporal evolution of the Rényi-2 mutual information (49) of two modes ($l=\epsilon$) separated by a distance $d=(M+1)\epsilon$ after quenching the optical vacuum (45) to the fractional Laplacian theory (34) with $L=2,m=1,\epsilon=0.1$ fixed. We vary the coupling range [see inset in (a) for the coupling strength $f_{M}=\lvert f_{jj+M}(\alpha)\rvert$ of the two modes at various distances $M$] from $\alpha=2$ (relativistic) in (b) to $\alpha\to 0$ (all-to-all) in (f). We cut off the mutual information below $0.01$ to probe the first correlation front (points) by fitting the characteristic curves summarized in Tab. 1 and shown in (a). The light cones start to bend when $\alpha<2$, as long-range couplings enable distant modes to become correlated at a faster rate. Correlations between neighboring modes pass the threshold at later times for longer range couplings, see gray regions in (b)(f), since the nearest-neighbor coupling strength $f_{0}$ decreases monotonically with $\alpha$, see inset in (a), converging to $f_{M}\to 1/2$ when $\alpha\to 0$.
V.2.2 Correlation spreading
We now study how two equally sized spatial regions of length $\ell$ separated by a distance $d>0$ become correlated after the quench, as shown in the inset of Fig. 4(b). To this end, we consider the Gaussian Rényi-$2$ mutual information
$\begin{split}I_{2}(t,\ell,d)&=2S_{2}(t,\ell)-S_{2}(t,\ell\cup\ell)\\ &=\dfrac{1}{2}\ln\frac{\det\sigma^{2}(t,\ell)}{\det\sigma(t,\ell\cup\ell)},\end{split}$ (49)
where $S_{2}(t,\ell\cup\ell)$ and $\sigma(t,\ell\cup\ell)$ are associated with the combined region $[0,\ell]\cup[\ell+d,2\ell+d]$. Analogously to the Rényi-$2$ entanglement entropy, the latter equals the Wigner mutual information for Gaussian states [127, 128, 129].
In the quasi-particle picture and the infinite volume limit, the relativistic QFT (29) predicts
$\begin{split}I_{2}(t,\ell,d)&=\int\dfrac{\mathrm{d}p}{2\pi}s_{2}(p)\Big{[}\max\left(2|v(p)|t,d\right)\\ &\hskip 62.59596pt+\max\left(2|v(p)|t,d+2\ell\right)\\ &\hskip 62.59596pt-2\max\left(2|v(p)|t,d+\ell\right)\Big{]},\end{split}$ (50)
see App. D.1 for finite-size corrections.
We benchmark the photonic simulation in Fig. 4(b) for $d=L/5$, with all other parameters left unchanged. The two spatial regions become correlated only after the light cones of the fastest quasi-particles intersect, i.e., for $t\geq\tau$. We remark that this is the essence of Lieb-Robinson bounds [133, 30, 134, 135]. Further, correlations decrease around the revival time $5\tau$, with the dip’s width being broadened compared to the Rényi-2 entanglement entropy as a result of having more possibilities for the quasi-particle configuration to resemble the initial state.
Strikingly, the early period $t\lesssim 3\tau$ is well captured by simulations with small system sizes $N\sim 5-10$. We find deviations from the QFT prediction around the revival time. Here, reasonable agreement requires around $N\sim 20$ modes, which is still accessible on existing platforms.
V.3 Systematic study: Bending of light cones
Finally, we investigate the rate at which two individual modes ($l=\epsilon$) at a distance $d=(M+1)\epsilon$ become correlated as the coupling range is varied. The onset of correlations is expected to follow (curved) light-cones as predicted by Lieb-Robinson bounds, which have been recently extended to long-range theories in the context of spin systems characterized by finite-dimensional Hilbert spaces [134] (see also [135]). Substantially fewer strict results are available for bosonic systems and the quasi-particle picture aside from numerical studies [102, 103, 104].
Type Form Coupling range
Linear $t\gtrsim d$ $\alpha=2$
Algebraic $t\gtrsim d^{\gamma},\,\gamma\in(0,1)$ $21\leq\alpha<$ 2
Logarithmic $t\gtrsim\log d$ $10<\alpha<$ 1
Constant $t\gtrsim\text{const}.$ $\alpha\to 0$
Table 1: Functional forms of the light cones describing the onset of correlations between two modes separated by a distance $d$ for varying coupling range $\alpha\in(0,2]$. The found regimes agree with those predicted in [134, 135].
By exploiting the OTA’s configurability, we simulate the fractional Laplacian theory (34) with $\alpha\in(0,2]$ and probe the build-up of correlations in space and time with the Rényi-2 mutual information following (49), see Fig. 5, where $L=2,m=1,\epsilon=0.1$ such that $N=20$ as in the previous example. Depending on the range $\alpha$, we observe the light cones to follow characteristic curves, see Fig. 5(a) and Tab. 1 for an overview.
As discussed in Sec. V.2, correlations follow straight light cones in a relativistic theory ($\alpha=2$), with the incline given by the fastest quasi-particle velocity, see Fig. 5(b). This is caused by the coupling strength $f_{M}=\lvert f_{jj+M}(\alpha)\rvert$ between two modes separated by $M$ modes [see inset in Fig. 5(a)] being non-zero only if the modes are adjacent, i.e., $f_{M}\propto\delta_{M0}$ when $\alpha=2$. Local correlations between neighboring modes require a finite time to reach the threshold $I_{2}=0.01$, as indicated by the gray region.
In the regime $21\leq\alpha<$ 2, see Figs. 5(c)-(d), the light cones start to bend, demonstrating faster correlation propagation for large distances, as expected for long-range couplings characterized by $f_{M}>0$ for all $M$. More precisely, the light cones follow algebraic curves. Simultaneously, nearest-neighbor correlations appear delayed, which is caused by a decreasing coupling strength between neighboring modes, as shown in the upper curve of the inset in Fig. 5(a). Upon further increasing the coupling range to $10<\alpha<$ 1, the light cones become logarithmic, see Fig. 5(e), until they disappear in the limit $\alpha\to 0$, see Fig. 5(f). The latter case corresponds to an all-to-all coupling of the modes with $f_{M}\to 1/2$ for all distances $M$, allowing for immediate information transfer and rendering all modes pari passu.
Remarkably, our findings are in perfect agreement with recent results for spin systems [134, 135] at a moderate system size of $N=20$ modes. We stress that only two of the twenty modes have to be measured and that the maximum required squeezing decreases approximately linearly with coupling range $\alpha$, with the maximum occurring at $\alpha=2$ for which $\lvert z_{\text{max}}\rvert\approx 0.35\approx 3$dB. This highlights the prospects of the OTA for systematically engineering and studying the dynamics of quantum information in various field-theoretic settings beyond the simple case of a relativistic scalar.
VI Discussion
We established the Optical Time Algorithm (OTA) for quantum-simulating the dynamics of various free scalar field theories. Our encoding is both intuitive and efficient: a single layer of phase shifters evolves the eigenmodes in time, while the physical properties of the underlying theory are encoded in two fixed beam splitter arrays followed by single-mode squeezers. In this way, a wide landscape of QFT dynamics can be quantum-simulated by adapting the parameters of a few optical elements, including (real- and complex-valued) relativistic, non-local, and non-relativistic theories as well as curved spacetimes (also exotic combinations of the latter can be envisioned), with little to no constraints on the theories’ parameters, the simulation time, or the spacetime dimension. Regarding applications, we focused on the light-cone-like spreading of quantum entanglement and correlations after a quench for varying coupling range, in which case the OTA reduces to a squeezing layer followed by passive linear optical elements. We found that $10-20$ modes are sufficient to capture field-theoretic predictions, while only a handful of those modes need to be measured.
We envision four main routes for generalizations and further applications. First, in analyzing the information dynamics of quantum fields, we have focused on pure Gaussian initial states such as the vacuum, in which case the OTA attains an elementary form. This was mainly motivated by the ubiquity of Gaussian states in quantum optics experiments. However, the OTA does not rely on simple input states. Given the growing range of techniques for preparing non-Gaussian states in the lab, for instance, via conditioning schemes [136, 137], the OTA is readily applicable for studying non-Gaussian scenarios. In this context, it is of particular interest to consider the conditioning of the particle numbers at (some of) the output modes using photon-number-resolving detectors, for which the experimental setup is precisely given by the Gaussian boson sampler. This would grant access to the symmetry sectors of the (local) particle number operator, which, given a suitable readout technique, could enable the readout of symmetry-resolved entanglement entropies [20, 138].
Second, while the OTA maintains its form in time-dependent scenarios whenever the diagonalization transformation (realized by the BS layers) remains time-independent, the situation becomes substantially more complex when considering general time dependencies, which are evident, for example, in the context of cosmology [27] and driven Bose-Einstein condensates [20, 42]. In this case, acting with the unitary time-evolution operator on the quadratures necessitates the inclusion of time-ordering into the exponential of the corresponding symplectic matrix transformation. This induces time dependencies in the symplectic eigenvalues, thereby complicating the disentangling of temporal contributions in this more general setting.
Third, as the OTA covers all relevant free theories, a natural extension would be the incorporation of interactions. The experimental realization of strong interactions at the level of few photons is challenging since nonlinear optical effects in conventional materials are typically negligible at the associated light powers. However, in the last two decades significant advances have been made towards this goal [139], e.g., by interfacing optical modes with single atoms or atomic ensembles in cavity QED systems [140] and by using hybrid states of light and atomic media [141, 142, 143]. We speculate that the key idea of our approach, that is, isolating the time-dependence and encoding the Hamiltonian into a fixed optical circuit, is generalizable to weakly interacting theories in a perturbative fashion. Beyond the perturbative regime, a Trotterization of the dynamics might be necessary.
Fourth, our application of the OTA for studying information dynamics constitutes a proof of principle regarding the accessibility of continuum physics with a modest number of optical modes and moderate squeezing. For this and other potential applications, a more in-depth analysis of experimental constraints, such as noise and photon losses, is warranted. To this end, one may consider alternative decomposition schemes for the beam splitter layer, which can be more robust against experimental imperfections [144, 145].
Recent advances towards scalable multimode photonic platforms using spatial [49, 146, 147, 148, 149], temporal [51, 150], and frequency  [151] degrees of freedom of light provide exciting opportunities for experimental realizations of the quantum field simulations proposed in this work. Although we emphasized photonic implementations due to their high level of configurability, other setups concerned with bosonic modes can also be envisioned, including circuit QED [152] and ultracold atoms [97] trapped in highly configurable optical potentials [18, 23].
Acknowledgements
We thank René Sondenheimer for useful discussions. T.H. acknowledges support from the European Union under project ShoQC within the ERA-NET Cofund in Quantum Technologies (QuantERA) program, as well as from the F.R.S.- FNRS under project CHEQS within the Excellence of Science (EOS) program. M.D. and M.G. are supported by funding from the German Research Foundation (DFG) under the project identifier 398816777-SFB 1375 (NOA).
Data availability and source code
We provide the code used for generating Figs. 4 and 5 together with an algorithm that computes all circuit parameters given a Hamiltonian matrix $H$ via the GitHub repository [90].
Appendix A Proofs
A.1 Proof of Eq. (15)
The isolation of the time parameter into single-mode phase shifters is achieved by generalizing Williamson’s theorem for the Hamiltonian matrix $H=S_{\text{SD}}^{\intercal}DS_{\text{SD}}$ [see Eq. (8)] to the symplectic matrix $S(t)$ describing time evolution. After introducing the corresponding unitary transformation $\bm{U}_{\text{SD}}$ defined via $S_{\text{SD}}\bm{r}=\bm{U}_{\text{SD}}^{\dagger}\bm{r}\bm{U}_{\text{SD}}$, the Hamiltonian becomes
$\bm{H}=\dfrac{1}{2}\bm{r}^{\intercal}H\bm{r}=\dfrac{1}{2}\bm{U}_{\text{SD}}^{\dagger}\bm{r}^{\intercal}(D_{N}\oplus D_{N})\bm{r}\bm{U}_{\text{SD}}.$ (51)
By linearity, the corresponding unitary can be rewritten as a product of time-dependent phase shifts between two time-independent unitaries [cf. Eq. (11)]
$\bm{U}(t)=e^{-it\bm{H}}=\bm{U}_{\text{SD}}^{\dagger}\bm{U}_{\text{PS}}(\varphi(t))\bm{U}_{\text{SD}},$ (52)
where $\varphi(t)=t(d_{1},\ldots,d_{N})^{\intercal}$. At the level of symplectic matrices, the previous decomposition reads
$S(t)=S_{\text{SD}}^{-1}S_{\text{PS}}(\varphi(t))S_{\text{SD}}.$ (53)
A.2 Proof of the OTA
A.2.1 Eigensystems of $H^{\phi}$ and $H^{\pi}$
For the OTA, we consider the class of block diagonal Hamiltonians with commuting blocks $H^{\phi}$ and $H^{\pi}$. Given Eq. (53), our goal is to further decompose the SD matrix by means of optical transformations. We make the ansatz
$S_{\text{SD}}=\begin{pmatrix}\mathds{O}_{N}&A_{N}\\ B_{N}&\mathds{O}_{N}\end{pmatrix},$ (54)
with $A_{N}$ and $B_{N}$ being further constrained as $S_{\text{SD}}$ must diagonalize the Hamiltonian matrix $H$ and has to be symplectic. When inserting (54) into (8), the former condition translates into the two constraints
$A_{N}^{\intercal}D_{N}A_{N}=H^{\pi},\quad B_{N}^{\intercal}D_{N}B_{N}=H^{\phi}.$ (55)
The latter condition can be evaluated by demanding that (54) preserves the symplectic form $\Omega$, which leads to
$A_{N}^{\intercal}B_{N}=-\mathds{1}_{N}.$ (56)
The constraints ensure $B^{\intercal}_{N}$ to be the (real) eigenvector matrix diagonalizing both $H^{\phi}$ and $H^{\pi}$. Then, multiplying the first two constraints and using the third to remove $A_{N}$, results in
$D^{2}_{N}=(B_{N}^{\intercal})^{-1}H^{\phi}H^{\pi}B_{N}^{\intercal},$ (57)
which implies Eq. (19). Further, the constraints fix the normalization of $B_{N}B_{N}^{\intercal}$: isolating $B_{N}B_{N}^{\intercal}=B_{N}H^{\phi}B_{N}^{-1}D_{N}^{-1}$ from the second constraint and using the similarity relation for $H^{\phi}$ on the right hand side, which itself can be rewritten in terms of $B_{N}^{-1}$ since $H^{\phi}$ is symmetric, leads to
$B_{N}B_{N}^{\intercal}=\text{diag}(\lambda^{\phi}_{1}/d_{1},\dots,\lambda^{\phi}_{N}/d_{N})\equiv\Gamma_{N}.$ (58)
By using Eq. (19), it becomes evident that the elements of the diagonal matrix $\Gamma_{N}=\text{diag}(\gamma_{1},\dots,\gamma_{N})$ are related to the eigenvalues of $H^{\phi}$ and $H^{\pi}$ via Eq. (18). We remark that the same line of reasoning applies when working with $A_{N}$ instead of $B_{N}$.
A.2.2 Squeezer layer and gauge fixing
Since the normal decomposition is not unique [88, 89], the SD matrix (53) is defined only up to an orthosmyplectic transformation [^3] At the level of $H$, this is equivalent to $G^{\intercal}DG=D$. $S_{\text{SD}}\to GS_{\text{SD}}$, with $G\in OrSp(2N)$ leaving the phase shifter array invariant, i.e.,
$G^{-1}S_{\text{PS}}(\varphi(t))G=S_{\text{PS}}(\varphi(t)).$ (59)
We refer to this mathematical redundancy encapsulated by $G$ as a gauge transformation.
The freedom of choosing the gauge comes in handy when decomposing $S_{\text{SD}}$ into optical elements, for which we employ the (symplectic) polar decomposition [70]
$S_{\text{SD}}=\Sigma P.$ (60)
Here, we introduced the symplectic, symmetric, and positive semi-definite matrix $\Sigma=(S_{\text{SD}}S_{\text{SD}}^{\intercal})^{1/2}$, while $P=\Sigma^{-1}S_{\text{SD}}$ is orthosymplectic. After evaluating the matrix $\Sigma^{2}$ for (54) in some general gauge $G$, we find
$\Sigma^{2}=\begin{pmatrix}G_{2}\Gamma_{N}G_{2}^{\intercal}+G_{1}\Gamma^{-1}_{N}G_{1}^{\intercal}&G_{2}\Gamma_{N}G_{1}^{\intercal}-G_{1}\Gamma^{-1}_{N}G_{2}^{\intercal}\\ G_{1}\Gamma_{N}G_{2}^{\intercal}-G_{2}\Gamma^{-1}_{N}G_{1}^{\intercal}&G_{1}\Gamma_{N}G_{1}^{\intercal}+G_{2}\Gamma^{-1}_{N}G_{2}^{\intercal}\end{pmatrix},$ (61)
where we used (56).
The latter matrix decouples when choosing the gauge $G_{1}=\mathds{O}_{N}$ and $G_{2}=\mathds{1}_{N}$, in which case
$\Sigma=\Gamma_{N}^{1/2}\oplus\Gamma_{N}^{-1/2}=S_{\text{Sq}}(z),$ (62)
represents a squeezer with $z_{j}=-(1/2)\ln\gamma_{j}$ [cf. Eq. (13)], as desired. Further, the remaining matrix $P$ indeed encodes the orthonormal eigenvectors of $H^{\phi}$ (or $H^{\pi}$) since
$P=(\Gamma_{N}^{-1/2}B_{N})\oplus(\Gamma_{N}^{-1/2}B_{N})\equiv P_{N}\oplus P_{N},$ (63)
where we recall that $\Gamma_{N}$ is diagonal. Note also that the symplectic constraint (56) becomes $B_{N}=-\Gamma_{N}A_{N}$, confirming that $A_{N}$ and $B_{N}$ are proportional.
We remark that our choice of gauge fixing improves the efficiency over a direct implementation of $S_{\text{SD}}$ as in (54): choosing $G_{1}=\mathds{1}_{N}$ and $G_{2}=\mathds{O}_{N}$ (corresponding to $G=\mathds{1}_{2N}$) leads to $P_{N}$ appearing on the anti-diagonal blocks of $P$ instead. This would require an additional layer of phase shifters – which amounts to rotating $P$ back to diagonal form – before the beam splitter array.
A.2.3 Decomposing the beam splitter array
Inspired by [66], we further decompose the orthosymplectic matrix $P_{N}$ by means of a QR decomposition
$P_{N}O_{N}=\mathcal{P}_{N},$ (64)
where $\mathcal{P}_{N}=\text{diag}(\pm 1,\dots,\pm 1)$ and $O_{N}$ being an orthogonal matrix, which can be written as a matrix product over Givens rotations $O_{jk}(\theta)$ defined as
$O_{jk}(\theta)=\begin{pmatrix}1&\cdots&0&\ldots&0&\cdots&0\\ \vdots&\ddots&\vdots&&\vdots&&\vdots\\ 0&\cdots&\cos(\theta)&\cdots&\sin(\theta)&\cdots&0\\ \vdots&&\vdots&\ddots&\vdots&&\vdots\\ 0&\cdots&-\sin(\theta)&\cdots&\cos(\theta)&\cdots&0\\ \vdots&&\vdots&&\vdots&\ddots&\vdots\\ 0&\cdots&0&\ldots&0&\cdots&1\\ \end{pmatrix}.$ (65)
The latter contains the beam splitter transformation $S_{\text{BS}}(\theta_{jk})$, i.e., a rotation, at the $j$th and $k$th intersections between rows and columns.
We engineer a sequence of Givens rotations that results in a diagonal matrix where all but the last entry are one, i.e., $\mathcal{P}_{N}^{\pm}=\text{diag}(1,\dots,\pm 1)$. Denoting by $p_{jk}$ the elements of the initial matrix $P_{N}$, we choose for all $k$ such that $1\leq k<N$, as long as there is a column $j>k$ with $p_{kj}$ different from zero, the beam splitter transmittivity $\theta$ and some $r>0$ such that
$r\cos\theta=p_{kk},\quad-r\sin\theta=p_{kj}.$ (66)
When applying the Givens rotation implicitly defined via the latter equations to $P_{N}$ from the right $P_{N}\to P_{N}O_{jk}(\theta)$, the corresponding elements of the transformed matrix become $p_{kk}=r$ and $p_{kj}=0$. Repeating this procedure for all $1\leq k<N$ and $j>k$ indeed results in
$P_{N}O_{12}O_{13}\dots O_{1N}O_{23}\dots O_{N-1N}=\mathcal{P}^{\pm}_{N},$ (67)
which is equivalent to
$P=(\mathcal{P}_{N}^{\pm}\oplus\mathcal{P}_{N}^{\pm})S^{-1}_{\text{BS}}(\theta),$ (68)
with Eq. (14) understood. In this way, all beam splitter parameters can be identified iteratively.
The last step is to treat the two possible signs of $\mathcal{P}_{N}^{\pm}$ in Eq. (68). When the last entry is positive, we obtain $P=S^{-1}_{\text{BS}}(\theta)$, which is what we wanted to show. For the negative sign, we require an additional phase shift on the last mode $P=S_{\text{PS}}(0,\dots,\pi)S^{-1}_{\text{BS}}(\theta)$. However, this phase can be absorbed by adjusting the gauge: using that $S_{\text{PS}}(0)=S_{\text{PS}}(\pi)=-\mathds{1}_{2}$ is proportional to the identity, we can rewrite the full symplectic diagonalizing matrix as $S_{\text{SD}}=S_{\text{Sq}}(z)P=S_{\text{PS}}(0,\dots,\pi)S_{\text{Sq}}(z)S^{-1}_{\text{BS}}(\theta)$, where we used the standard gauge fixing ($G_{1}=\mathcal{O}_{N},G_{2}=\mathds{1}_{N}$). Transforming the gauge to $G\to S_{\text{PS}}^{-1}(0,\dots,\pi)G$, such that $G_{2}=\mathcal{P}_{N}^{-}$, supersedes the additional phase shift, while leaving the squeezer unchanged. It is straightforward to check that this gauge still satisfies (59), which completes our proof.
Appendix B Circuit parameters for the relativistic scalar on $N=5$ modes
To illustrate the decomposition scheme, we provide all circuit parameters for the relativistic theory (30) when $N=5$. The phases $d_{j}$ reduce to the relativistic dispersion (32), while the squeezing parameters are $z_{j}=-(1/2)\ln(\epsilon d_{j})$. The beam splitter parameters $\theta_{ij}$ follow after a straightforward numeric computation using the iterative scheme described in App. A.2.3. We report all the parameters for $m=1$ and $\epsilon=2$ in Tab. 2.
$\theta_{ij}$ $j=1$ $2$ $3$ $4$ $5$
$i=1$ $2.36$ $0.62$ $0.52$ $0.46$
$2$ $0.52$ $-0.62$ $-0.79$
$3$ $-1.02$ 00
$4$ $-1.57$
$5$
$d_{j}$ $1$ $1.16$ $1.38$ $1.38$ $1.16$
$z_{j}$ $-0.35$ $-0.42$ $-0.51$ $-0.51$ $-0.42$
Table 2: Optical parameters for the relativistic theory on $N=5$ modes with $m=1$ and $\epsilon=2$.
Appendix C Fractional Laplacian theory on a finite lattice
In flat space, the Laplace operator $-\partial_{x}^{2}$ is diagonalized by a Fourier transform to momentum space. Hence, arbitrary powers of the Laplace operator $(-\partial_{x}^{2})^{\alpha/2}$ can be treated by raising its momentum-space eigenvalues to the same power and then transforming back to position space [106]. The eigenvalue of the term $\phi(-\partial^{2}_{x})^{\alpha/2}\phi$ is dictated by the fractional dispersion relation (36) in the massless limit $m\to 0$. Adapting this logic to the periodic lattice results in
$f_{jj^{\prime}}(\alpha)=\dfrac{2^{\alpha}}{N}\sum_{k\in\mathcal{K}}e^{-2\pi i(j-j^{\prime})k/N}\left\lvert\sin^{\alpha}\left[\tfrac{\pi k}{N}\right]\right\rvert,$ (69)
with the DFT being defined as
$\tilde{\phi}_{k}=\dfrac{1}{\sqrt{N}}\sum_{j\in\mathcal{J}}e^{-2\pi ijk/N}\phi_{j},$ (70)
and discrete momenta labeled by $k\in\mathcal{K}=\{0,\dots,N-1\}$.
Appendix D Quantum quench of the relativistic theory
D.1 Finite size effects
In a finite volume $L<\infty$ with periodic boundary conditions, quasi-particles have discrete momenta and reenter from the left (right) when traveling to the right (left) and passing through the boundary at $L$ (00). The integral over momenta in (48) is replaced by a sum and the condition $\min(2|v(p)|t,\ell)$, which can be expanded into the two terms $2|v(p)|t\times\theta(l-2|v(p)|t)+\ell\times\Theta(2|v(p)|t-l)$, has to be modified to account for reentrances of quasi-particles, see Fig. 6. Then, the Rényi-2 entanglement entropy becomes [82, 83]
$\begin{split}S_{2}(t,\ell)&=\sum_{g_{k}<\ell/L}s_{2,k}\,g_{k}+\dfrac{\ell}{L}\sum_{\ell/L\leq g_{k}<1-\ell/L}s_{2,k}\\ &\hskip 11.38092pt+\sum_{g_{k}\geq 1-\ell/L}s_{2,k}(1-g_{k}),\end{split}$ (71)
where we sum over the momentum modes $k$ and introduced the function $g_{k}=\{2|v_{k}|t/L\}$, with $\{x\}$ denoting the fractional part of $x$.
A similar argument can be made for adapting the Rényi-2 mutual information (49) to finite system sizes. Due to the circular geometry, the two subregions are separated by $d$ as well as the complementary distance $d^{\prime}=L-2\ell-d$ as quasi-particles contributing to the mutual information can also move in the opposite direction, yielding one additional contribution, see left vs. right panel in Fig. 6. Further, each quasi-particle pair contributes to the mutual information twice: once after having traveled a distance $(L+d)/2$, but also for $(L+d^{\prime})/2$, see the green diamonds in Fig. 6. In total, there are four contributions to mutual information before revivals, i.e., this process repeating itself. Hence, after introducing the function
Figure 6: Trajectories of quasi-particles emerging at $A_{1}$ and $A_{2}$ and moving towards each other directly (left panel) or indirectly (right panel). This results in one (three) additional finite-size contribution(s) in the Rényi-2 entanglement entropy (71) [mutual information (73)] compared to the infinite volume formulas presented in the main text. The ballistic motion is periodic, leading to momentum-dependent revivals.
$\begin{split}i_{2,k}(x)=s_{2,k}\Big{[}&\max\left(x/L,g_{k}\right)\\ &+\max\left(x/L+2\ell/L,g_{k}\right)\\ &-2\max\left(x/L+\ell/L,g_{k}\right)\Big{]},\end{split}$ (72)
the finite-size mutual information takes the form
$\begin{split}I_{2}(t,\ell,d)=\sum_{k\in\mathcal{K}}\Big{[}&i_{2,k}(d)+i_{2,k}(d^{\prime})\\ &+i_{2,k}(L+d)+i_{2,k}(L+d^{\prime})\Big{]}.\end{split}$ (73)
D.2 Entropy density of the GGE and mode populations
The entropy density $s_{2,k}$ follows from computing the Rényi-2 entanglement entropy (47) of the GGE $\bm{\rho}(\infty,l)\propto\text{exp}(-\sum_{k}y_{k}\bm{n}_{k})$ in the eigenbasis of the quasi-particle number operator, leading to [81, 82]
$S_{2}(\infty,\ell)=S_{2,\text{GGE}}=\sum_{k\in\mathcal{K}}s_{2,k},$ (74)
with coefficients
$s_{2,k}=\ln\left[n_{k}^{2}-(1-n_{k})^{2}\right],$ (75)
where $n_{k}=\braket{0}{\bm{n}_{k}}{0}$ specifies the quasi-particle content of the initial state (that is conserved during time evolution).
To compute $n_{k}$, we first determine the eigenmodes of the relativistic theory (30). As discussed in Sec. IV.2, performing a DFT (70) to momentum space renders the corresponding Hamiltonian diagonal, to wit
$\bm{H}_{\text{R}}=\sum_{k\in\mathcal{K}}\omega_{k}\left(\tilde{\bm{a}}_{k}^{\dagger}\tilde{\bm{a}}_{k}+\dfrac{1}{2}\right).$ (76)
Therein, we introduced the mode operators
$\begin{split}\tilde{\bm{a}}_{k}&=\dfrac{1}{\sqrt{2\epsilon\omega_{k}}}\left(\epsilon\omega_{k}\tilde{\bm{\phi}}_{k}+i\tilde{\bm{\pi}}_{k}\right),\\ \tilde{\bm{a}}_{k}^{\dagger}&=\dfrac{1}{\sqrt{2\epsilon\omega_{k}}}\left(\epsilon\omega_{k}\tilde{\bm{\phi}}_{k}^{\dagger}-i\tilde{\bm{\pi}}_{k}^{\dagger}\right),\end{split}$ (77)
describing quasi-particle excitations. Note here that the appearance of the relativistic dispersion $\omega_{k}$ guarantees Lorentz-invariance of the momentum integral measure in the continuum.
The ground state |$\ket{\Omega}$ of (30), or, equivalently, (76), corresponds to the quasi-particle vacuum $\tilde{\bm{a}}_{k}\ket{\Omega}=0$, while the input-state vacuum $\bm{a}_{j}\ket{0}=0$ contains $n_{k}$ quasi-particles. The latter follows after expanding the quasi-particle mode operators $(\tilde{\bm{a}}_{k},\tilde{\bm{a}}_{k}^{\dagger})$ in terms of the position-space, i.e., photonic, mode operators $(\bm{a}_{j},\bm{a}_{j}^{\dagger})$, for which we combine Eqs. (77) and (70) with $\bm{a}_{j}=(\bm{\phi}_{j}+\bm{\pi}_{j})/\sqrt{2}$, leading to
$\tilde{\bm{a}}_{k}=\dfrac{1}{2\sqrt{N}}\sum_{j\in\mathcal{J}}e^{-2\pi ijk/N}(\eta_{k}\bm{a}_{j}+\kappa_{k}\bm{a}_{j}^{\dagger}),$ (78)
with
$\eta_{k}=\dfrac{1}{\sqrt{\epsilon\omega_{k}}}+\sqrt{\epsilon\omega_{k}},\quad\kappa_{k}=\dfrac{1}{\sqrt{\epsilon\omega_{k}}}-\sqrt{\epsilon\omega_{k}}.$ (79)
This way, we find
$\begin{split}n_{k}&=\braket{0}{\tilde{\bm{a}}_{k}^{\dagger}\tilde{\bm{a}}_{k}}{0}\\ &=\dfrac{\kappa_{k}^{2}}{4N}\sum_{j,j^{\prime}\in\mathcal{J}}e^{-2\pi i(j-j^{\prime})k/N}\langle 0|\bm{a}_{j^{\prime}}\bm{a}_{j}^{\dagger}|0\rangle\\ &=\dfrac{\kappa_{k}^{2}}{4}\\ &=\dfrac{1}{4}\left(\dfrac{1}{\epsilon\omega_{k}}+\epsilon\omega_{k}-2\right).\end{split}$ (80)
D.3 Comment on divergences in the continuum limit
The input state chosen in the main text, which is experimentally motivated, leads to divergences in the continuum. Formally, the corresponding pre-quench Hamiltonian (45) attains the form
$\bm{H}_{0}=\dfrac{1}{2}\int\mathrm{d}x\left[\bm{\pi}^{2}+(\lim_{\epsilon\rightarrow 0}\epsilon^{-2})\bm{\phi}^{2}\right]$ (81)
in the continuum. It becomes apparent that this Hamiltonian has a divergent mass term. This causes further divergences in the corresponding momentum modes’ energy ($\omega_{0,k}=1/\epsilon$), population numbers ($n_{k}\sim 1/\epsilon$), and thus also the entropy density $s_{2,k}\sim-\ln\epsilon$. We remark that, as of working within a regularized theory, these divergences can be cured by renormalizing the considered quantities with appropriate functions of $\epsilon$.